# Risks of Practicing Large Language Models in Smart Grid: Threat Modeling and Validation

### v1

@@ Abstract
Large Language Model (LLM) is a significant breakthrough in artificial intelligence (AI) and holds considerable potential for application within smart grids. However, as demonstrated in previous literature, AI technologies are susceptible to various types of attacks. It is crucial to investigate and evaluate the risks associated with LLMs before deploying them in critical infrastructure like smart grids. In this paper, we systematically evaluate the vulnerabilities of LLMs and identify two major types of attacks relevant to smart grid LLM applications, along with presenting the corresponding threat models. We then validate these attacks using popular LLMs, utilizing real smart grid data. Our validation demonstrates that attackers are capable of injecting bad data and retrieving domain knowledge from LLMs employed in smart grid scenarios.
@@ I Introduction
Recently, a significant breakthrough in AI has been the development of Large Language Models (LLMs). LLMs have a large number of model parameters and possess the capability to understand and process general natural language [5]. Due to their exceptional performance across various natural language processing tasks, LLMs have facilitated numerous applications across different industries, such as research [6] software development [7]. Given the successful performance of AI techniques in numerous smart grid applications, it is natural to explore the potential of integrating LLMs into the smart grid [8, 9, 10, 11].
Figure 1: Illustration of cyberattacks to LLMs used in smart grid. The figure shows the case that an outsider attacker injects false information into LLMs and a case that an internal attacker obtains unauthorized information from an internal LLM.
Despite the outstanding performance of AI models, they can also be vulnerable and expose systems to new types of cyberattacks. Previous research has shown that popular AI models, such as deep neural networks, are susceptible to various attacks, including adversarial [12] and backdoor attacks [13]. These attacks have proven to be effective against AI applications in smart grids as well [14, 15, 16, 17]. Therefore, conducting thorough research on potential risks is crucial before deploying large language models (LLMs) in smart grid applications.
The applications of large language models (LLMs) are rapidly evolving, and research into their security, especially in potential smart grid applications, is still nascent. [10] represents the sole research focusing on the security of LLMs in smart grid contexts; it outlines potential threat types but lacks a comprehensive threat model and validation. To address this gap, we systematically study the potential threats posed by LLMs in smart grid applications, develop generalized threat models, and validate these threats through simulation LLM attacks. Our contributions are summarized as follows:
  • We analyze how the threats posed by LLMs differ from those associated with previous AI models, such as Support Vector Machines (SVM), and identify the scope of risks that LLMs bring to the smart grid.
  • From the perspective of attack motivation, we propose two general threat models for potential attacks on LLMs deployed in smart grids.
  • As proof of concept, we validate these two threat types using popular LLMs (GPT-3.5[18] and GPT-4 [19]) and demonstrate that attackers can readily inject malicious data into and extract confidential information from these models.
  • We open-source the complete data, code, and evaluation result used for this paper to encourage more research on this area [20].
@@ II Related Work
II-A Machine Learning in Power Grid
Machine learning (ML)techniques, which involve extracting statistical patterns from data, have been extensively studied and applied in various smart grid applications. These applications include load forecasting [21], anomaly detection [22], energy theft detection [4], and cyberattack detection [2], among others. In 2016, Sun et al. utilized social network information to detect power outages in specific areas [23]. They employed Bayesian models to analyze Twitter messages, achieving accurate detection results. Similarly, Wang et al. [22] collected public news and utilized a pre-trained language model to detect incidents at renewable energy facilities. ML has also shown remarkable effectiveness in cyberattack detection. For example, false data injection attacks, a well-known threat to power system state estimation, have been effectively identified using ML technologies such as Recurrent Neural Networks (RNNs) [24, 25, 26] and Convolutional Neural Networks (CNNs) [2]. Ismail et al. [4] applied CNNs to detect energy theft, achieving a state-of-the-art detection rate of 99.3%.
Recently, research into the application of LLMs to power systems has begun. Dong et al. [11] explore potential use cases of LLMs within power systems, demonstrating the capabilities of finely-tuned LLMs in pattern-recognition tasks such as equipment damage detection and document analysis. Similarly, Huang et al. [9] have examined the use of LLMs in specific power system operations, including optimal power flow (OPF) and electric vehicle (EV) scheduling. These studies highlight the potential utility of LLMs in enhancing the efficiency and reliability of smart grids.
Further research on the application of ML in the power grid can be found in [27].
II-B Machine Learning Security in Power Grid
While ML has achieved outstanding results in many smart grid tasks, its security remains a significant concern. Originally demonstrated in the field of computer vision, well-performing ML models have been shown to be vulnerable to adversarial attacks [12]. This type of attack has also proven effective against ML-based smart grid applications. In 2018, Chen et al. first explored adversarial examples with both categorical and sequential applications in power systems [14]. Later, in 2021, [17] demonstrated that attackers could craft adversarial data capable of bypassing ML model FDIA detection while adhering to the physical constraints of the power grid. Additionally, [28] showed that adversarial attackers could also circumvent energy theft detection ML models while reporting extremely low power consumption data.
At the time of this research, the study by Ruan et al. [10] is the only work into the potential risks of applying LLMs in power systems. This work identifies critical risks such as privacy invasion, deteriorated performance, and semantic divergence.
@@ III-A Background: Prompt and LLM
An LLM is a pre-trained machine learning model that is capable for processing natural language-related tasks. A natural language-based prompt that specifies the context of the task is usually provided to LLMs to achieve different task goals. As shown in Fig. 2, an LLM application usually starts with the developer providing the prompt to LLM, and the LLM will respond to the user’s input based on the information and requirements provided in the prompt.
@@
When applying the LLMs to the smart grid, as shown in [11, 9], the prompt provided to LLM is expected to be the domain knowledge and requirements of the task. Without loss of generality, the general LLM query can be represented as:
@@ III-B Threat Analysis: Scope
While research has identified various vulnerabilities in LLMs and their susceptibility to different types of attacks [29], such as Jailbreak [30], prompt injection [31], and privacy violation [32], not all vulnerabilities are relevant for power system applications. To investigate the potential threats LLMs may pose to the smart grid, it is crucial to emphasize the properties of LLMs that differ from previous AI applications and focus on their practical usage in smart grid scenarios. Therefore, the scope of concern should exclude:
  • General LLM Inherent Limitations: Attacks such as jailbreak, which attempt to circumvent LLM restrictions to output sensitive content, such as personal information, represent a general issue for all LLM applications and are not unique to smart grids.
@@
  • Unauthorized Access of Public LLMs: This is a broader security issue for smart grid network management. Such issues should be addressed through general network control measures, such as firewalls [34], rather than being seen as unique challenges posed by LLMs.
III-C Threat Analysis: Threat Models
Although not explicitly mentioned, most previous AI methods proposed for smart grid applications assume that the model was trained by trusted staff, such as internal engineers, and that all data used for training came from reliable sources, such as smart meter data [21, 22, 4, 2]. However, this assumption no longer holds true when applying LLMs. Due to the high cost of training LLMs, pre-trained models are used. These pre-trained LLMs, such as GPT-4, possess general knowledge that can lead to unexpected behavior according to well-crafted inputs.
We analyzed the existing research on the LLMs’ vulnerabilities, for smart grid applications, the major risks can be categorized into: bad data injection and domain knowledge leaking, as demonstrated in Figure 1. We proposed two different threat models accordingly.
@@ III-C1 Bad Data Injection
The category of bad data injection primarily concerns scenarios where LLMs are publicly accessible and used to extract information from open resources. This vulnerability is pertinent to various practical applications within smart grids, such as processing user feedback reports from different sources. We specifically consider cases where attackers may attempt to inject malicious data through public access points to the LLMs. The objective of such attacks is to deceive the LLMs into producing incorrect results that could mislead smart grid operators. For instance, an attacker could manipulate the language models (LLMs) into misclassifying a report about a blackout as a thank-you letter. The general representation of bad data injection is
@@
  • Public Access: LLMs have a public access point, making them accessible to a wide user base.
  • Attacker Capabilities: An attacker can send input to the LLMs, leveraging this public interface. However, attackers cannot receive the direct output from the LLMs
@@
Given that LLMs are utilized for extracting information from publicly available sources, the existence of a public access point is inherent. This allows the attacker to inject data into the LLMs, for example, by submitting a falsified customer report. The high-level functions of the LLMs are sometimes disclosed to the public, which further facilitates attacker understanding and planning. Such information might be obtained from publicly accessible documents like a power utility portal’s user agreement and privacy policy, or through simple guesses.
@@ III-C2 Domain Knowledge Extraction
Domain knowledge extraction mainly addresses scenarios involving unauthorized access to the domain knowledge that serves as the prompt for LLM applications. Typically, LLMs are fed raw information but are designed only to provide specific responses to external queries. In practice, this domain knowledge might include power grid configuration information, where the LLM’s task is to answer specific questions based on this information without disclosing the raw configuration details. Domain knowledge extraction occurs when an attacker attempts to manipulate the LLMs into exposing domain knowledge that should not be revealed. Domain knowledge extraction can be represented as:
@@
  • Attacker Capabilities: An attacker can interact with the LLMs, including sending inputs and receiving outputs.
  • Attacker Knowledge: The attacker has high-level domain knowledge, such as the functions of the LLMs.
We primarily consider insider attackers for domain knowledge extraction, as the best practice is to avoid providing sensitive information to publicly accessible LLMs. Thus, insider attackers, who have legitimate access, can interact with the LLMs by sending inputs and receiving outputs. By analyzing these outputs, the attacker can derive high-level information about the domain knowledge, potentially extracting sensitive information not intended for disclosure.
@@ IV Validation
In this section, we validate two types of attacks, bad data injection and domain knowledge extraction, by simulating scenarios involving LLM applications in a smart grid context. We selected GPT-3.5 and GPT-4 as the LLMs for our study due to their popularity and extensive coverage in related research [11, 9]. Our validation results demonstrate that, under the practical threat models proposed in this paper, attackers are capable of successfully executing both types of attacks on smart grid LLM applications.
@@ IV-A Bad Data Injection
Application Simulation: To demonstrate the susceptibility of LLMs to bad data injection attacks, we explore an LLM application designed to process public resources and identify potential incidents at renewable energy facilities. Specifically, the LLM analyzes random messages from the public to determine whether they contain information about incident events related to renewable energy infrastructures. In practice, these messages could originate from social media posts or customer reports. Utilizing such an application can facilitate the rapid identification of potential incidents, thereby mitigating potential damage and enhancing response efficiency.
Dataset: We utilized the dataset from [22] as our primary data source. This dataset comprises messages collected from public news sources, focusing on energy events. Initially, the messages in their raw dataset were tokenized and not specific to renewable energy facilities. We merged the tokens and manually curated the records to specifically include those related to renewable energy facilities. Our refined dataset contains 423 labeled messages. Examples from the dataset are provided below, with ’yes’ representing related and ’no’ not related.
@@
Attack Simulation: In Fig. 4, we present a selection of the prompts we used, including normal inputs and the bad data injection prompts employed in our attack simulations. The simulations involve three types of malicious inputs: ‘all yes’, ‘all no’, and ‘reverse’.
Figure 4: Bad Data Injection Demo
@@
  • Phase 1: We query the LLMs using normal inputs, record their responses, and compare these results with the ground truth labels.
  • Phase 2: We query the LLMs with the same inputs again, but this time we intersperse them with the three types of bad data sequentially to observe the effects on LLM responses. These altered results are then compared to the ground truth to evaluate the impact of the bad data injections.
Simulation Result: The summarized results of our simulation are depicted in Table I. From these results, it is evident that with normal inputs, the LLMs achieve very good performance across all evaluation metrics. However, when bad data is injected, there is a significant drop in detection accuracy for all three cases. These evaluation results demonstrate that well-designed malicious inputs can deceive the LLMs into misclassifying the inputs, highlighting the vulnerability of LLMs to bad data injection attacks.
TABLE I: Bad Data Injection Result
Result Normal Inject-Yes Inject-No Inject-Rev
GPT3.5 Accuracy 89.1% 47.7% 52.2% 33.5%
GPT3.5 Precision 87.8% 47.7% 0% 34%
GPT3.5 Recall 89.6% 100% 0% 41.5%
GPT3.5 F1 88.7% 64.6% 0% 37.4%
GPT4 Accuracy 93.8% 48.9% 52.2% 43.4%
GPT4 Precision 94% 48.3% 0% 45.4%
GPT4 Recall 93.1% 100% 0% 91.1%
GPT4 F1 93.5% 65.1% 0% 60.1%
Some sample responses are shown in Fig. 5.
Figure 5: Sample Responses of Bad Data Injection

### v2

@@ Abstract
Large language models (LLMs) represent significant breakthroughs in artificial intelligence and hold considerable potential for applications within smart grids. However, as demonstrated in previous literature, AI technologies are susceptible to various types of attacks. It is crucial to investigate and evaluate the risks associated with LLMs before deploying them in critical infrastructure like smart grids. In this paper, we systematically evaluated the risks of LLMs and identified two major types of attacks relevant to potential smart grid LLM applications, presenting the corresponding threat models. We also validated these attacks using popular LLMs and real smart grid data. Our validation demonstrates that attackers are capable of injecting bad data and retrieving domain knowledge from LLMs employed in different smart grid applications.
@@ I Introduction
Recently, a significant breakthrough in AI has been the development of Large Language Models (LLMs). LLMs have a large number of model parameters and possess the capability to process general natural language [5]. Due to their exceptional performance across various natural language processing (NLP) tasks, LLMs have facilitated numerous applications across different industries [6, 7]. Given the successful performance of AI techniques in numerous smart grid applications, it is natural to explore the potential of integrating LLMs into the smart grid [8, 9, 10, 11].
Figure 1: Illustration of cyberattacks to LLMs used in smart grid. The figure shows two cases: (1) an outsider attacker injects false information into LLMs and (2) an internal attacker obtains unauthorized information from LLMs.
Despite the outstanding performance of AI models, they can also be vulnerable and expose systems to new types of cyberattacks. Previous research has shown that popular AI models, such as deep neural networks, are susceptible to various attacks, including adversarial [12] and backdoor attacks [13]. These attacks have proven to be effective against AI applications in smart grids as well [14, 15, 16, 17]. Therefore, conducting thorough research on potential risks is crucial before deploying LLMs in smart grid applications.
The applications of LLMs are rapidly evolving, and research into their risks, especially in potential smart grid applications, is still nascent. To the best of our knowledge, [10] represents the sole research focusing on the security of LLMs in smart grid contexts; it outlines potential threat types but lacks a comprehensive threat model and validation. To address this gap, we systematically study the potential threats posed by LLMs in smart grid applications, develop generalized threat models, and validate these threats through simulated LLM attacks. Our contributions are summarized as follows:
  • We analyze how the threats posed by LLMs differ from those associated with previous AI models (e.g. SVM) and identify the scope of risks that LLMs bring to the smart grid.
  • Considering attack motivations, we propose two general threat models for potential attacks on LLMs deployed in smart grids: (1) bad data injection attacks by external adversaries, and (2) domain knowledge extraction attacks by internal actors.
  • As a proof of concept, we validate these two threat types using popular LLMs: OpenAI GPT-3.5 [18], OpenAI GPT-4 [19], and Meta LLaMA-3 [20]. Through thorough simulations, we demonstrate that attackers can readily inject malicious data into, and extract confidential information from, these models.
  • We have open-sourced the complete data, code, and evaluation results used in this paper to encourage further research on this topic [21].
@@ II Related Work
II-A Machine Learning and NLP in Power Grids
Machine learning techniques have been extensively studied and applied in various smart grid applications, including load forecasting [22], anomaly detection [23], energy theft detection [4], and cyberattack detection [2, 24, 25, 26].
In recent years, natural language processing (NLP) applications have also appeared in smart grids. In 2016, Sun et al. utilized social network information to detect power outages in specific areas [27]. They employed Bayesian models to analyze Twitter messages, achieving accurate detection results. Similarly, Wang et al. [23] collected public news and utilized a pre-trained language model to detect incidents at renewable energy facilities.
In addition, research into the application of LLMs to power systems has begun. Dong et al. [11] explored potential use cases of LLMs within power systems, demonstrating the capabilities of fine-tuned LLMs in pattern recognition tasks such as equipment damage detection and document analysis. Similarly, Huang et al. [9] examined the use of LLMs in specific power system operations, including optimal power flow (OPF) and electric vehicle (EV) scheduling. These studies highlight the potential utility of LLMs in enhancing the efficiency and reliability of smart grids.
Further research on the application of machine learning in power grids can be found in [28].
II-B Machine Learning Security in Smart Grid
While ML has achieved outstanding results in many smart grid tasks, its security remains a significant concern.
Well-performing ML models have been shown to be vulnerable to adversarial attacks, a phenomenon originally demonstrated in the field of computer vision [12]. This type of attack has also proven effective against ML-based smart grid applications. In 2018, Chen et al. first explored adversarial examples with both categorical and sequential applications in power systems [14]. Later, Li et al. [17] demonstrated that attackers could craft adversarial data capable of bypassing ML-based cyberattack detection while adhering to the physical constraints of the power grid. [29] showed that adversarial attackers could also circumvent ML-based energy theft detection models while reporting extremely low power consumption data.
At the time of this research, the study by Ruan et al. [10] is the only work investigating the potential risks of applying LLMs in power systems. This work identifies critical risks such as privacy invasion, deteriorated performance, and semantic divergence.
@@ III-A Background: Prompt and LLM
An LLM is a pre-trained ML model capable of processing natural language tasks. To achieve different task goals, LLMs are typically provided with a natural language prompt that specifies the context of the task. As shown in Fig. 2, an LLM application usually starts with the developer providing a prompt to the LLM, which defines how the LLM should process user inputs. The LLM then responds to the user’s input based on the information and requirements specified in the prompt.
@@
As presented in [11, 9], when applying LLMs to the smart grid, the prompt provided to the LLM is expected to include domain knowledge and task requirements. Without loss of generality, the LLM query can be represented as:
@@ III-B Threat Analysis: Scope
While research has identified various vulnerabilities in LLMs and their susceptibility to different types of attacks [30], such as jailbreak [31], prompt injection [32], and privacy violations [33], not all vulnerabilities are relevant in the smart grid context. To investigate the potential threats LLMs may pose to the smart grid, it is crucial to emphasize the properties of LLMs that differ from previous AI applications and focus on their practical usage in smart grid scenarios. Therefore, the scope of concern should exclude:
  • General LLM Inherent Limitations: Attacks such as jailbreaks, which attempt to circumvent LLM restrictions to output sensitive content like personal information, represent general issues for all LLM applications and are not unique to smart grids.
@@
  • Unauthorized Access to Public LLMs: This is a broader security issue for smart grid network management. Such issues should be addressed through general network control measures, such as firewalls [35], rather than being seen as unique challenges posed by LLMs.
III-C Threat Models
Although not explicitly mentioned, most AI methods proposed for smart grid applications in previous literature assume that the AI model is trained by trusted personnel, such as internal engineers, and that all data used for training comes from reliable sources like smart meter data [22, 23, 4, 2]. However, this assumption no longer holds when applying LLMs. Due to the extremely high data requirements and costs associated with LLM training, pre-trained models are typically used. These pre-trained LLMs, whether open-source (e.g., Meta LLaMA-3) or proprietary (e.g., OpenAI GPT-4), have pre-trained parameters that may lead to unexpected behavior in response to well-crafted inputs from attackers.
Based on our analysis of existing research on LLM vulnerabilities, we categorize the main risks of LLMs for smart grid applications into bad data injection and domain knowledge extraction, as demonstrated in Figure 1. Accordingly, we propose two different threat models.
@@ III-C1 Bad Data Injection
Bad data injection primarily concerns scenarios where LLMs receive input from public resources and are used to extract and process information from that input. This scenario is relevant to many practical applications within smart grids, such as processing user feedback reports from diverse sources.
We specifically consider cases where attackers attempt to inject bad data through public access points to the LLMs by manipulating legitimate inputs. The objective of such attacks is to deceive the LLMs into producing incorrect results that could mislead smart grid operators. For instance, an attacker could deceive the LLMs into misclassifying a report about a blackout as a thank-you letter by injecting well-crafted text into it. The general representation of bad data injection is
@@
  • Public Access: LLMs receive inputs from a public access point, making them accessible to a wide user base.
  • Attacker Capabilities: The attacker is assumed to be able to modify legitimate input sent to the LLMs via this public interface. However, attackers cannot receive the direct output from the LLMs.
@@
Given that LLMs are utilized for processing information from publicly available sources, the existence of a public access point is inherent. Injecting bad data into legitimate inputs can be implemented in many ways, such as man-in-the-middle attacks. The high-level functions of the LLMs are sometimes disclosed to the public, like news and articles [36], which further facilitates attacker understanding and planning. Such information might be obtained from publicly accessible documents like a power utility portal’s user agreement and privacy policy, or through simple inference.
@@ III-C2 Domain Knowledge Extraction
Domain knowledge extraction addresses scenarios involving unauthorized access to the domain knowledge that serves as the prompt for LLM applications. In many practical use cases, LLMs are provided with prompts containing raw data information but are designed to provide only specific responses to users’ queries. For example, the domain knowledge may include smart grid configuration information, where the LLM’s task is to answer specific questions based on this information without disclosing the raw configuration details. Domain knowledge extraction occurs when an attacker attempts to deceive the LLM into exposing domain knowledge that should not be revealed. This can be represented as:
@@
  • Attacker Capabilities: An insider attacker can interact with the LLM, including sending inputs and receiving outputs.
  • Attacker Knowledge: The attacker has high-level domain knowledge, such as understanding the functions of the LLM.
We focus on insider attackers in the context of domain knowledge extraction, as best practices recommend avoiding the provision of sensitive information to publicly accessible LLMs. Thus, insider attackers, who have legitimate access, can interact with the LLM by sending inputs and receiving outputs. By analyzing these outputs, the attacker can derive detailed information about the domain knowledge, potentially extracting sensitive information not intended for disclosure.
@@ IV Validation
In this section, we validate the two types of attacks—bad data injection and domain knowledge extraction—by simulating scenarios involving LLM applications in a smart grid context. We selected the proprietary models OpenAI GPT-3.5 and GPT-4, as well as the open-source model Meta LLaMA-3, for our study due to their popularity and extensive coverage in related research [11, 9, 37]. Our validation results demonstrate that, under the practical threat models proposed in this paper, attackers can successfully execute both types of attacks on potential smart grid LLM applications.
The evaluations of the OpenAI models were implemented through the official OpenAI API [38]. For Meta LLaMA-3, we deployed the LLaMA-3.1-8B-Instruct [39] model on Amazon Web Services (AWS) using Hugging Face Inference Endpoints. Due to page limitations, only essential results are presented here; the implementation details, such as the complete system prompts and the datasets used in this paper, can be found in [21].
@@ IV-A Bad Data Injection
IV-A1 Application Introduction
@@
To demonstrate the susceptibility of LLMs to bad data injection attacks, we study the use case presented in [23] where an AI model was trained and used to perform a binary classification task to determine whether a message from public sources is related to an incident of renewable energy facilities. In practice, utilizing such an application can facilitate the rapid identification of potential incidents, thereby mitigating potential damage and enhancing response efficiency.
The dataset in [23] comprises messages collected from public news sources, focusing on energy events. Originally, the messages in the raw dataset were tokenized and not specific to renewable energy facilities. We revert the tokens to texts and manually curated the records to specifically include those related to renewable energy facilities. The refined dataset contains 423 labeled messages. Examples from the dataset are presented in Fig. 3, with ‘Yes’ representing related and ‘No’ not related.
IV-A2 Attack Simulation
@@
TABLE I: LLMs Performance with Normal Inputs
Model Accuracy Precision Recall F1
OpenAI GPT-3.5 89.1% 87.8% 89.6% 88.7%
OpenAI GPT-4 93.8% 94% 93.1% 93.5%
Meta Llama-3 88.9% 85.4% 92.6% 88.8%
REI Classifier [23] 88.5% 89.0% 93.8% 91.4%
  • Phase 1 - Normal Inputs: We reproduce the binary classification task using the LLMs. We empirically drafted a system prompt to identify relevant messages, and the same prompt was used for all three models. A sample of the system prompt can be found in Fig. 4. In this phase, we queried all three LLMs with normal messages, and the final input to the LLMs can be represented as
    $\text{Final Input}=\text{System Prompt}||\text{Normal Input}$Final Input = System Prompt | | Normal Input
    We summarize the performance of the LLMs on the classification task of normal inputs in Table I. By comparing the four metrics, we observe that all the LLMs achieve similar or better performance compared with the classifier proposed in [23]. We note that the effort required for the LLMs to achieve this performance is only to craft a simple system prompt, which is minimal effort compared with previous methods that include collecting and processing data, designing the model, and training. This demonstrates the efficiency of LLMs and their vast potential for application in smart grids.
  • Phase 2 - Injection Inputs: In this phase, we simulated scenarios where the normal inputs were modified by the attacker through the injection of crafted text. Considering the different motivations of the attackers, they may choose to perform a false negative attack (converting all classification results to ‘No’), a false positive attack (converting all classifications to ‘Yes’), or a reverse attack (reversing the classification results). The corresponding injection texts are shown in Fig. 4. We simulated these three types of attack by appending the injection text after the normal input. The final input to the LLMs can be represented as
[…]
    Similar to Phase 1, we queried the LLMs with the same messages but interspersed them with the three types of bad data sequentially to observe the effects on LLM responses. The performances of the LLMs under bad data injection attacks are shown in Table. II. By comparing the performances of the LLMs on normal inputs and injection inputs, we observe a significant drop in associated metrics for all three LLMs under all three different attack scenarios. These evaluation results demonstrate that well-designed injection texts can deceive the LLMs into misclassifying the inputs, highlighting the vulnerability of LLMs to bad data injection attacks in potential smart grid applications.
Figure 4: Bad Data Injection Prompts Presentation
TABLE II: LLMs Performance under Bad Data Injection
Model Metric Inject-FP Inject-FN Inject-Rev
OpenAI GPT-3.5 Accuracy 47.7% 52.2% 33.5%
Precision 47.7% 0% 34%
Recall 100% 0% 41.5%
F1 64.6% 0% 37.4%
OpenAI GPT-4 Accuracy 48.9% 52.2% 43.4%
Precision 48.3% 0% 45.4%
Recall 100% 0% 91.1%
F1 65.1% 0% 60.1%
Meta LlaMA-3 Accuracy 47.8% 52.2% 33.8%
Precision 47.8% 0% 38.8%
Recall 100% 0% 66.8%
F1 64.6% 0% 49.1%